A single die is tossed thirteen times. Find the probability that exactly eleven of the tosses show a "three."
Answer by edjones(7311)

The chance of a 3 on one toss is 1/6
Let a=5/6 and b=1/6
(a+b)^13
a^2b^11 is the term we need.
the coefficient of this term is nCr=13C11=78
78a^2b^11
=78*(5/6)^2*(1/6)^11
=325/2176782336 or 1.49303*10^-7 A very small number, indeed.
.
Ed